Tell what will happen to each standard error of the mean when the sample size:


1. increased from 50 to 200? Explain


2. decreased from 250 to 225? Explain


3. increased from 200 to 400? Explain


4. increased from 40 to 50? Explain


5. decreased from 600 to 40? Explain



1
Expert's answer
2021-11-22T14:58:11-0500

"SE=\frac{s}{\sqrt{n}}."

So,

1) SE decreases by "\sqrt{\frac{200}{50}}=2" times.

2) SE increases by "\sqrt{\frac{250}{225}}\approx 1.0541" times.

3) SE decreases by "\sqrt{\frac{400}{200}}=1.4142" times.

4) SE decreases by "\sqrt{\frac{50}{40}}\approx 1.1180" times.

5) SE increases by "\sqrt{\frac{600}{40}}\approx 3.8730" times.